OpenStudy (anonymous):

Find the area of the region enclosed by the curve y = 1/sqrt(x) and x-axis on (0,1]. Is this convergent or divergent?

OpenStudy (ujjwal):

You should integrate 1/sqrt(x) from 0 to 1 to find the area! \[\int_{0}^{1} \frac{1}{\sqrt x}=2\] So, area should be 2 sq units!

OpenStudy (anonymous):

What Is -x+4y-3z=11

OpenStudy (anonymous):

Thank you, ujjwal. And this is convergent because 2 is constant, right?

OpenStudy (ujjwal):

I have no idea about convergent and divergent are... So, sorry i can't help you with that!

OpenStudy (anonymous):

That's fine. I found the answer. Thank you though

OpenStudy (ujjwal):

WC :)